Marc is 8 years younger than his brother Richard. The sum of their ages is 22 years. What equation could be used to find Marc's age? What is Marc's age?
Answer by stanbon(75887)

Marc is 8 years younger than his brother Richard. The sum of their ages is 22 years. What equation could be used to find Marc's age? What is Marc's age?
--------
Equations:
m = r - 8
m + r = 22
----
Rearrange:
m-r = 8
M+r = 22
-------
Add and solve for "m":
2m = 30
Marc's age is 15 yrs.
------
M+r = 22
r = 7 yrs
=============
Cheers,
Stan H.
=========